OpenStudy (anonymous):

Use the first derivative rule to find the values of x for which f(x) = 1/3 x^3 + 3/2 x^2 - 18 x + 3 is increasing and decreasing. Hint: Use what you know about the shape of cubic polynomial graphs from Algebra and Precalculus. . The function f is increasing for the following values of x: it should look something like a

OpenStudy (anonymous):

here's another one lol

OpenStudy (cwrw238):

yea! f(x) = 1/3 x^3 + 3/2 x^2 - 18 x + 3 f'(x) = x^2 + 3x - 18 when f is increasing x^2 + 3x - 18 > 0 critical points (x + 6)(x - 3) = 0 x = -6 and x = 3 f increasing when x < -6 when x is in between -6 and 3 x is decreasing when x is > 3 its increasing hmm - i'm getting a bit confused here - ill draw a rough graph

OpenStudy (cwrw238):